Joe's Taxi Service charges a flat rate plus an additional $0.45 per mile. If someone takes one of Joe's taxis for 13 miles and is charged $12.35, determine the
following:
a: What does Joe charge for the flat rate?
b: Write an equation that represents the total cost of riding in Joe's taxi as a function of the number of miles driven.
c: If someone rides in one of Joe's taxis and pays $23.15, how many miles did he travel?

Answer by Boreal(15235)

C=FR+0.45m
$12.35=FR+0.45(13); the last term=$5.85
$6.50=Flat Rate.
-
$23.15=FR+0.45m=$6.50+0.45m
$16.65=0.45m
m=37 miles